Accountancy

KleerView [KW] sells several high quality electronic products. One of these is an electronic screen

which it sells at $1,700 each with a standard one-year warranty. It also sells to its customers a separate

three-year extended warranty, commencing after the end of the standard warranty. KW offers its

customers who have purchased such extended warranties a service, when necessary, to perform either

appropriate repairs or to replace the defective unit.

The company estimates, based upon its experience from prior years, the total warranty costs for the

standard warranty to average $50 per screen, being $30 for parts and $20 for labor. It also expects the

average three-year extended warranty costs to be $80 for parts and $160 for labor for each contract. It

further assumes that the warranty costs for the extended warranty contracts will be incurred as follows:

20% in 20X2, 50% in 20X3 and the balance in 20X4. And finally, KW expects to recognize warranty

revenues based on the proportion of costs incurred out of the total estimated costs.

During 20X1, the company sold 600 screens and 540 extended warranty contracts for cash. During the

year, it also incurred some actual costs associated with the standard warranties related to the 20X1 sales

of screens. The cost for parts were 150% of the labor costs. On December 31, 20X1, KW reported the

following:

Current Liabilities:

Estimated Liability Under Warranty $ 15,300

Unearned Warranty Revenue $151,200

KW incurred actual costs associated with the standard warranties related to the 20X1 sales of screens, in

20X2 amounting to $17,700 and the cost for parts therein amounted to $11,800. It further incurred actual

costs associated with the extended warranty contracts which were consistent with what the company had

expected to incur.

Required:

1. Determine the selling price for each extended warranty contract.

2. Present all journal entries to be prepared, in proper format, in 20X1 in order to record all of the

warranty related transactions and adjustments for 20X1.

3. Present all journal entries to be prepared, in proper format, in 20X2 in order to record all of the

transactions related to the standard warranties of 20X1.

4. Present all journal entries to be prepared, in proper format, in 20X2 in order to record all of the

transactions related to the extended warranties of 20X1.

5. What liabilities related to warranties, would be reported on the December 31, 20X2 Balance Sheet.

Show how these would be classified.
Answer #1

1) No income is booked for standard warranty expense because it comes with product so unearned warranty income is fully related to the extended warranty contract Selling price per extended warranty contract = Unearned warranty revenue / Extended warranty contracts \(\Rightarrow 151,200 / 540\)

\(=>280\)

2) Estimated standard warranty cost \(=600 * 50=>30,000\) Estimated warranty liability in future \(=15,300\) Warranty expense paid \(=30,000-15,300=>14,700\)

3) Warranty expense in \(20 \mathrm{X} 2\) Warranty expense incurred \(=17,700+11,800=>29,500\) Warranty liability payable balance at the beginning of the year \(=15,300\) Additional warranty expense incurred \(=29,500-15,300=>14,200\) Debit Warranty expense 14,200

Debit Warranty liability payable 15,300 Credit cash 29,500

4) At the end of \(1^{\text {st }}\) year entity has estimated the warranty liability to be 15,300 and the amount of cash paid is 14,700 out of warranty expense estimated. Units sold during the year may have occurred at the beginning, end or during the year so amount paid refers to the warranty period expired units in the year itself and balance left has an expiry period after an year or it may be units sold at the end, hence for warranty period expired units revenue and expense has to be booked based on remaining years.

Warranty expired units in \(20 \mathrm{X} 1=\) Warranty Cost / Cost per unit \(=>14,700 / 50\)

\(=>294\) units Warranty expired units in \(20 \mathrm{X} 2=600-294=>306\) units Out of total 600 units; customer of 540 units have availed extended warranty; so taking proportionately number of units having additional warranty units that have expired can be found \(=294 * 540 / 600=>265\) units round off Additional Warranty units not expired at the end of \(20 \mathrm{X} 1=540-260=>175\) units Warranty income per year \(=(280 * 265\) units \() / 3=24,733\) Income to be booked for \(20 \mathrm{X} 2=265\) units \(* 240(160+80) / 3 \Rightarrow 21,200\) Journal entry Debit Unearned warranty income \(\quad 24,733\) Credit Warranty income 24,733

Debit Warranty expense

21,200

Credit Warranty expense payable 21,200

5) Warranty liabilities that are expected to be incurred in next 12 months are treated as current liabilities so warranty liability payable created has to be recorded as current liability for 21,200

Unearned warranty income Unearned warranty income which is expected to be earned in next 12 months have to be treated as current liability and balance amount has to be treated as long term liability Unearned warranty income at the end of \(20 \mathrm{X} 2=151,200-24,733=>126,467\) Unearned warranty income to be reversed in next 12 months \(=540 * 280 / 3=>50,400\) Unearned income to be treated as current liability \(=50,400\) Unearned income to be treated as long term liability \(=126,467-50,400=>76,067\)

answered by: whoami